OpenStudy (anonymous):

okay so using Binomial Theorem determine the 5th term of (x+2b)^4 without fully expanding.

OpenStudy (anonymous):

find the 5th term of \[ {(x+2b)}^4\] without fully expanding?

OpenStudy (athe):

\[C_4^4(x)^{0}(2b)^4=16b^4\]

OpenStudy (amistre64):

thats a nice way of getting around the nCr binom latex stuff

OpenStudy (athe):

use this \[(a+b)^n=\sum_{k=0}^{n}C_n^k a^{n-k}b^k\] there \[C_n^k=\frac{n!}{(n-k)!k!}, k!=k \times (k-1)\times (k-2)\dots 2\times1\]

OpenStudy (amistre64):

^nC_r works too i think

OpenStudy (amistre64):

\[^n C_r\]

OpenStudy (anonymous):

I like \( \binom{n}{r} \) Just to add @amistre64 lol